Find an equation of the line that is parallel to the line y=-4x+5 and that passes through the point(-3,-2)
Answer by mananth(12270)

y=-4x+5 and that passes through the point(-3,-2)
..
slope of the line = -4
parallel line will have same slope = -4
point (-3-2)
..
Y = MX +B
TO FIND B
plug value of the point in the equation
-2 = -4*-3 +b
-2 = 12 +b
b= -14
The equation will be
slope = -4, b = -14
y = -4x -14